1.    The specific heat capacity of zinc is 0.388 J/g.˚C. Calculate the initial temperature of zinc block of mass 38g if the amount of heat supplied to it was 220 J and the maximum temperature reached by the zinc block was 55˚C.



1
Expert's answer
2022-02-01T09:08:02-0500

"H=mc\Delta\theta"


"220=38×0.388(55-\theta)"

"55-\theta=14.92"


"\theta=40.08°C"